Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

All envelope contains eight bills......

Expert replies
by pzazz12 » Fri Oct 22, 2010 3:54 am
All envelope contains eight bills:2 ones,2 fives ,2 tens,and 2 twenties.Two bills are drawn at random without replacement .what is the probability that their sum is $20 or more?

A.1/4
B.2/7
C.3/7
D.1/2
E.2/3
Join the discussion
Source: — Problem Solving |

by Rahul@gurome » Fri Oct 22, 2010 4:39 am
pzazz12 wrote:All envelope contains eight bills:2 ones,2 fives ,2 tens,and 2 twenties.Two bills are drawn at random without replacement .what is the probability that their sum is $20 or more?

A.1/4
B.2/7
C.3/7
D.1/2
E.2/3
Let us calculate the inverse probability.
If the sum of two bills is less than $20, the bills must be a pair from the set {$1, $1, $5, $5, $10, $10} with a reduced pair ($10, $10).

Total number of selection of 2 bills from 6 bills = 6C2 = 15.
Total number of selection of 2 bills from 8 bills = 8C2 = 28.
Probability that the sum of the selected two bills is less than $20 = (15-1)/28 = 1/2

Therefore, probability that the sum of the selected two bills is $20 or more= 1 - 1/2 = 1/2.

The correct answer is D.
Rahul Lakhani
Quant Expert
Gurome, Inc.
https://www.GuroMe.com
On MBA sabbatical (at ISB) for 2011-12 - will stay active as time permits
1-800-566-4043 (USA)
+91-99201 32411 (India)
Join the discussion

by abhishekg21 » Fri Oct 22, 2010 8:02 am
shouldn't it be 2/7.

total ways of selecting 2 cards out of 8 is =8*7=56 (considering the order)

out of options availbale(1,1,5,5,10,10,20,20) we can have 20 or more in following ways

(1,20),(5,20),(10,20),(20,20),(10,10)


now ways of selecting the above options considering the order
(1,20) in =2*2 =4
(5,20) =2*2=4
(10,20)=2*2=4
(20,20)=2*1=2
(10,10)=2*1=2
so prob is =16/56=2/7

please tell me what is wrong in this approach ?
Join the discussion

by Rahul@gurome » Fri Oct 22, 2010 8:25 am
Yes we can consider the order, but in that case also the result will be same. Because 2 will be multiplied both in the numerator and denominator.

Flaws in your procedure are in the second part.
abhishekg21 wrote: ...out of options availbale(1,1,5,5,10,10,20,20) we can have 20 or more in following ways

(1,20),(5,20),(10,20),(20,20),(10,10)

now ways of selecting the above options considering the order
(1,20) in =2*2 =4
(5,20) =2*2=4
(10,20)=2*2=4
(20,20)=2*1=2
(10,10)=2*1=2
so prob is =16/56=2/7

please tell me what is wrong in this approach ?
Considering the order (1,20), (5,20) and (10,20) can be selected in 8 ways not in 4 ways as there are 2 of each of them. Say for the pair (1,20), there are two $1 bill 1a and 1b, and two $20 bill 20a and 20b. Now considering the order, two of them can be selected in the following ways,
(1a,20a), (1a,20b), (1b,20a), (1b,20b), (20a,1a), (20a,1b), (20b,1a) and (20b,1b).

Therefore,
(1,20) in = 2*2*2 = 8 ways
(5,20) = 2*2*2 = 8 ways
(10,20)= 2*2*2 = 8 ways
(20,20)= 2*1 = 2 ways
(10,10)= 2*1 = 2 ways

So probability = (8+8+8+2+2)/56 = 28/56 = 1/2.
Rahul Lakhani
Quant Expert
Gurome, Inc.
https://www.GuroMe.com
On MBA sabbatical (at ISB) for 2011-12 - will stay active as time permits
1-800-566-4043 (USA)
+91-99201 32411 (India)
Join the discussion

by abhishekg21 » Fri Oct 22, 2010 8:34 am
thanks a lot.I got my mistake.
Join the discussion

by 5abi » Fri Oct 22, 2010 8:34 am
abhishekg21 wrote:shouldn't it be 2/7.

total ways of selecting 2 cards out of 8 is =8*7=56 (considering the order)

out of options availbale(1,1,5,5,10,10,20,20) we can have 20 or more in following ways

(1,20),(5,20),(10,20),(20,20),(10,10)


now ways of selecting the above options considering the order
(1,20) in =2*2 =4
(5,20) =2*2=4
(10,20)=2*2=4
(20,20)=2*1=2
(10,10)=2*1=2
so prob is =16/56=2/7

please tell me what is wrong in this approach ?
i got the same answer?
Join the discussion

by aspirant_mumbai » Fri Oct 22, 2010 10:45 pm
Hi All,

Here is my approach. I am using slot method.

We have two slots. Lets fix first one for 20, & the second slot for rest of 7 numbers.
So the total no of ways of selecting a sum greater than 20 is 1*7=7
Again repeating the same fixing 10 at slot one:
Total number of ways 1*3 (thats selecting 10,20,20) =3
Fixing 5 at slot one:
Total number of ways 1*2(selecting 20,20) = 2
For 1 at slot number 1
Total number of ways 1*2 (selecting 20,20) = 2

Total ways = 2*(7+3+2+2)=14*2

Total ways to selecting two numbers our of eight = 8*7
Probability = 14*2/8*7 = 1/2

Rahul Lakhani, please correct my approach if wrong.
Join the discussion

by Rahul@gurome » Fri Oct 22, 2010 11:36 pm
aspirant_mumbai wrote:Hi All,

Here is my approach. I am using slot method.

We have two slots. Lets fix first one for 20, & the second slot for rest of 7 numbers.
So the total no of ways of selecting a sum greater than 20 is 1*7=7
Again repeating the same fixing 10 at slot one:
Total number of ways 1*3 (thats selecting 10,20,20) =3
Fixing 5 at slot one:
Total number of ways 1*2(selecting 20,20) = 2
For 1 at slot number 1
Total number of ways 1*2 (selecting 20,20) = 2

Total ways = 2*(7+3+2+2)=14*2

Total ways to selecting two numbers our of eight = 8*7
Probability = 14*2/8*7 = 1/2

Rahul Lakhani, please correct my approach if wrong.
Your approach is fine. :)
Rahul Lakhani
Quant Expert
Gurome, Inc.
https://www.GuroMe.com
On MBA sabbatical (at ISB) for 2011-12 - will stay active as time permits
1-800-566-4043 (USA)
+91-99201 32411 (India)
Join the discussion

by aspirant_mumbai » Sat Oct 23, 2010 12:11 am
Thanks, Rahul :D
Join the discussion

by Scott@TargetTestPrep » Thu Jan 04, 2018 9:54 am
pzazz12 wrote:All envelope contains eight bills:2 ones,2 fives ,2 tens,and 2 twenties.Two bills are drawn at random without replacement .what is the probability that their sum is $20 or more?

A.1/4
B.2/7
C.3/7
D.1/2
E.2/3
The number of ways to select 2 bills from 8 bills is 8C2 = 8!/[2!(8-2)!] = (8 x 7)/2! = 28. Now we have to determine the number of ways to select 2 bills that total $20 or more.

If one of the two bills is one of the $20 bills, then the sum is guaranteed to be $20 or more. Thus, for each $20 bill selected, we can pair it with the 6 non-$20 bills. Since there are two $20 bills, we have 6 x 2 = 12 pairings. Furthermore, we have two more pairings: we can pair the two $20 bills with each other and we can also pair the two $10 bills with each other. Thus, we have 14 pairings of two bills that total $20 or more. Therefore, the probability is 14/28 = 1/2.

Answer: D

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion